# Timely Communications for Remote Inference

### v1

@@ Abstract
In this paper, we analyze the impact of data freshness on remote inference systems, where a pre-trained neural network infers a time-varying target (e.g., the locations of vehicles and pedestrians) based on features (e.g., video frames) observed at a sensing node (e.g., a camera). One might expect that the performance of a remote inference system degrades monotonically as the feature becomes stale. Using an information-theoretic analysis, we show that this is true if the feature and target data sequence can be closely approximated as a Markov chain, whereas it is not true if the data sequence is far from Markovian. Hence, the inference error is a function of Age of Information (AoI), where the function could be non-monotonic. To minimize the inference error in real-time, we propose a new “selection-from-buffer” model for sending the features, which is more general than the “generate-at-will” model used in earlier studies. In addition, we design low-complexity scheduling policies to improve inference performance. For single-source, single-channel systems, we provide an optimal scheduling policy. In multi-source, multi-channel systems, the scheduling problem becomes a multi-action restless multi-armed bandit problem. For this setting, we design a new scheduling policy by integrating Whittle index-based source selection and duality-based feature selection-from-buffer algorithms. This new scheduling policy is proven to be asymptotically optimal. These scheduling results hold for minimizing general AoI functions (monotonic or non-monotonic). Data-driven evaluations demonstrate the significant advantages of our proposed scheduling policies.
@@ I Introduction
A smaller AoI indicates more recent information at the receiver. While existing AoI research extensively analyzes and optimizes linear and nonlinear functions of AoI [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], there is a lack of comprehensive understanding on how to assess the value of fresh information in real-time systems. This gap motivates us to seek a more suitable analysis for evaluating the significance of fresh information for systems using it.
Previous studies [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] assumed system performance degrades monotonically with stale information (i.e., fresher data is always better). However, our remote inference experiments (see Sec. II-C) demonstrate that this assumption holds in some scenarios (e.g., video prediction), but not in others (e.g., temperature prediction or reaction prediction). For example, in predicting next hour temperature, temperature recorded at $24$ hours ago is more relevant than temperature recorded at $12$ hours ago due to daily weather patterns. Moreover, consider an example of reaction prediction: If a vehicle initiates braking, nearby vehicles don’t react instantly due to the response delay of the drivers or the braking systems. Therefore, slightly outdated actions can be more relevant for predicting reactions than the most current action. These observations highlight that fresher data is not always better. To accurately assess the value of fresh information, we need a robust analytical framework.
In this paper, we first present a new information-theoretic analysis to reveal when fresher data is always better and when it is not. In the second part of this paper, we focus on optimal communication system design for remote inference. Most existing studies in AoI literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] have focused on designing transmission scheduling strategies to minimize monotonic functions of AoI. However, the design of efficient scheduling policies for optimizing general, potentially non-monotonic functions of AoI remains unexplored. To that end, we design the first transmission scheduling policies to minimize general functions of AoI, regardless of whether the functions are monotonic or non-monotonic. The contributions of this paper are summarized as follows:
@@ I-A Related Works
The concept of Age of Information (AoI) has attracted significant research interest; see, e.g., [3, 5, 4, 20, 13, 21, 8, 7, 22, 23, 6, 9, 10, 11, 12, 14, 15, 16, 24, 25, 26, 27] and a recent survey [28]. Initially, research efforts were centered on analyzing and optimizing average AoI and peak AoI in communication networks [3, 5, 4, 20, 13]. Recent research endeavors on AoI have shifted towards optimizing the performance of real-time applications, such as remote estimation, remote inference, and control systems, by leveraging AoI as a tool. In [21, 8, 7, 23], information-theoretic metrics such as Shannon’s mutual information (or Shannon’s conditional entropy) has been used to quantify the amount of information carried by the received data about the current source value (or the amount of uncertainty regarding the current source value) as the data ages. In addition, a Shannon’s conditional entropy term $H_{\mathrm{Shannon}}(Y_{t}|X_{t-\Delta(t)}=x,\Delta(t)=\delta)$ was used in [22] to quantify information uncertainty given the most recent observation $X_{t-\Delta(t)}=x$ and AoI $\Delta(t)=\delta$. The information-theoretic metrics in these prior studies [21, 8, 7, 23, 22] cannot be directly used to evaluate system performance. To bridge the gap, in the present paper, we use an $L$-conditional entropy $H_{L}(Y_{t}|X_{t-\Delta(t)},\Delta(t))$, to approximate and analyze the inference error in remote inference, as well as the estimation error in remote estimation. For example, when the loss function $L(y,\hat{y})$ is chosen as a quadratic function $||y-\hat{y}||_{2}^{2}$, the $L$-conditional entropy $H_{L}(Y_{t}|X_{t-\Delta(t)},\Delta(t))=\mathbb{E}[(Y_{t}-\mathbb{E}[Y_{t}|X_{t-\Delta(t)},\Delta(t)])^{2}]$ is exactly the minimum mean squared estimation error in signal-agnostic remote estimation. This approach takes a significant step to bridge the gap between AoI metrics and real-world applications, by directly mapping the AoI to the application performance metrics.
In the earlier AoI studies [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], it was usually assumed that the observed data sequence is Markovian and the performance degradation caused by information aging was modeled as a monotonic AoI function. Hence, the earlier studies [4, 6, 5, 7, 12, 8, 9] adopted “generate-at-will” status updating model, where the transmitter can only select the most recently generated signal. However, practical data sequence may not be Markovian [7]. In the present paper, we propose a new local geometric approach to analyze both Markovian and non-Markovian time-series data. For non-Markovian time-series data, fresh data is not always better. To that end, we propose a new status updating model called the “selection-from-buffer” model, where the transmitter has the option to send any of the $B$ most recent features stored in the source buffer.
@@ II-B Offline Training and Online Inference
A supervised learning algorithm consists of two phases: offline training and online inference. In the offline training phase, a neural network based predictor is trained using one of the following two approaches.
@@
In the online inference phase, the pre-trained neural predictor is used to predict the target $Y_{t}$ in real-time. We assume that the process $\{(Y_{t},X_{t}),t=0,1,2,\ldots\}$ is stationary and the processes $\{(Y_{t},X_{t}),t=0,1,2,\ldots\}$ and $\{\Delta(t),t=0,1,2,\ldots\}$ are independent. Under these assumptions, if $\Delta(t)=\delta$, the inference error at time $t$ can be expressed as a function of the AoI value $\delta$, i.e.,
@@
where $P_{Y_{t},X_{t-\delta}}$ is the distribution of the target $Y_{t}$ and the feature $X_{t-\delta}$, and $\phi^{*}$ is the trained neural network. The proof of (4) is provided in Appendix E. In Sections V-VI, to minimize inference error, we will develop signal-agnostic transmission scheduling policies in which scheduling decisions are determined without using the knowledge of the signal value of the observed process. If the transmission schedule is signal-agnostic, then $\{(Y_{t},X_{t}),t=0,1,2,\ldots\}$ is independent of the AoI process $\{\Delta(t),t=0,1,2,\ldots\}$.
@@ II-C Experimental Results on Information Freshness
We conduct five remote inference experiments to examine how the training error and the inference error vary as the AoI increases. These experiments include (i) video prediction, (ii) robot state prediction in a leader-follower robotic system, (iii) actuator state prediction under mechanical response delay, (iv) temperature prediction, and (v) wireless channel state information prediction. In these experiments, we consider the quadratic loss function $L(\mathbf{y},\hat{\mathbf{y}})=\|\mathbf{y}-\hat{\mathbf{y}}\|^{2}_{2}$. Detailed settings of these experiments can be found in Appendix A. We present the experimental results of the first training method in Figs. 1-5. Related codes and datasets are accessible in our GitHub repository. [^1] https://github.com/Kamran0153/Impact-of-Data-Freshness-in-Learning To illustrate the training error of the second training method as a function of the AoI $\delta$, one can simply assess the training error using the training data samples with the AoI value $\delta$. As the results of the two training methods are similar, the experimental results of the second training method are omitted.
@@ III-A1 Training Error of the First Training Approach
$H_{L}(Y|X)=\min_{\phi(x)\in\mathcal{A},~{}\!\forall x\in\mathcal{X}}\mathbb{E}_{X,Y\sim P_{X,Y}}[L(Y,\phi(X))].$ (6)
@@ III-A2 Training Error of the Second Training Approach
$\!\!\!\!H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\Theta},\Theta)=\min_{\phi\in\Phi}\mathbb{E}_{Y,X,\Theta\sim P_{\tilde{Y}_{0},\tilde{X}_{-\Theta},\Theta}}\![L(Y,\!\phi(X,\Theta))],$ (10)
@@
which connects the information-theoretic lower bounds of $\mathrm{err_{training,1}}(\delta)$ and $\mathrm{err_{training,2}}(\delta)$.
@@ III-A3 Inference Error
$H_{L}(P_{Y};P_{\tilde{Y}}|P_{X})=\sum_{x\in\mathcal{X}}P_{X}(x)\mathbb{E}_{Y\sim P_{Y|X=x}}\!\!\left[L\!\left(Y,a_{P_{\tilde{Y}|\tilde{X}=x}}\right)\right]\!.$ (15)
If the function spaces $\Lambda$ and $\Phi$ are close to each other, the difference between $\mathrm{err_{inference}}(\delta)$ and the $L$-conditional cross entropy $H_{L}\left(P_{Y_{t}|X_{t-\delta}};P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}|P_{X_{t-\delta}}\right)$ is small.
@@
We first analyze the monotonocity of $L$-conditional entropy $H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\delta})$ as $\delta$ increases. If $\tilde{Y}_{0}\leftrightarrow\tilde{X}_{-\mu}\leftrightarrow\tilde{X}_{-\mu-\nu}$ is a Markov chain for all $\mu,\nu\geq 0$, by the data processing inequality for $L$-conditional entropy [35, Lemma 12.1], $H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\delta})$ is a non-decreasing function of $\delta$. Nevertheless, the experimental results in Figs. 1-5 show that the training error is a growing function of the AoI $\delta$ in some systems (see Fig. 1), whereas it is a non-monotonic function of $\delta$ in other systems (see Figs. 2-5). As we will explain below, a fundamental reason behind these phenomena is that practical time-series data for remote inference could be either Markovian or non-Markovian. For non-Markovian $(\tilde{Y}_{0},\tilde{X}_{-\mu},\tilde{X}_{-\mu-\nu})$, $H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\delta})$ is not necessarily monotonic in $\delta$.
@@
A Markov chain is an $\epsilon$-Markov chain with $\epsilon=0$. If $Z\rightarrow X\rightarrow Y$ is a Markov chain, then $Y\rightarrow X\rightarrow Z$ is also a Markov chain [42, p. 34]. A similar property holds for the $\epsilon$-Markov chain.
@@
Figure 5: Performance of channel state information (CSI) prediction. The training error and inference error are non-monotonic in AoI. As the feature sequence length $u$ increases, the errors tend closer to non-decreasing functions of the AoI.
@@
Using (8), (III-A1), and (15), it is easy to show that the $L$-conditional cross entropy $H_{L}(P_{Y_{t}|X_{t-\delta}};P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}|P_{X_{t-\delta}})$ is lower bounded by the $L$-conditional entropy $H_{L}(Y_{t}|X_{t-\delta})$. In addition, $H_{L}(P_{Y_{t}|X_{t-\delta}};P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}|P_{X_{t-\delta}})$ is close to its lower bound $H_{L}(Y_{t}|X_{t-\delta})$, if the conditional distributions $P_{Y_{t}|X_{t-\delta}}$ and $P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}$ are close to each other, as stated in Lemma 3.
@@
Given $\beta\geq 0$, if for all $\delta\in\mathbb{Z}^{+}$
@@ III-D Interpretation of the Experimental Results
We use Theorems 1-3 to interpret the experimental results in Figs. 1-5. In Fig. 1, the training and inference errors for video prediction are increasing functions of the AoI. This observation suggests that the target and feature time-series data $(Y_{t},X_{t-\mu},X_{t-\mu-\nu})$ for video prediction is close to a Markov chain. In the robot state prediction experiment depicted in Fig. 2, the state of the follower robot depends on the state of the leader robot received through a channel. Due to the communication delay from the leader robot to the follower robot, the target and feature data sequence $(Y_{t},X_{t-\mu},X_{t-\mu-\nu})$ can be far from a Markov chain. In the experiment of actuator state prediction under mechanical response delay, pole angle at time $t$ is strongly correlated with the cart velocity generated $25$ ms ago, as observed from data traces in Fig. 3(b). Moreover, temperature and CSI signals have long-range dependence. For example, temperature at time $t$ depends on the temperature of $24$ hours ago. These observations imply that the target and feature data sequence $(Y_{t},X_{t-\mu},X_{t-\mu-\nu})$ for all $\mu,\nu\geq 0$ may not be close to a Markov chain in the experimental results depicted in Figs. 2-5. Because the target and feature time-series data involved is non-Markovian, Theorems 1-3 suggest that the training error and inference error could be non-monotonic with respect to AoI, as observed in Figs. 2-5.

### v2

@@ Abstract
In this paper, we analyze the impact of data freshness on remote inference systems, where a pre-trained neural network infers a time-varying target (e.g., the locations of vehicles and pedestrians) based on features (e.g., video frames) observed at a sensing node (e.g., a camera). One might expect that the performance of a remote inference system degrades monotonically as the feature becomes stale. Using an information-theoretic analysis, we show that this is true if the feature and target data sequence can be closely approximated as a Markov chain, whereas it is not true if the data sequence is far from being Markovian. Hence, the inference error is a function of Age of Information (AoI), where the function could be non-monotonic. To minimize the inference error in real-time, we propose a new “selection-from-buffer” model for sending the features, which is more general than the “generate-at-will” model used in earlier studies. In addition, we design low-complexity scheduling policies to improve inference performance. For single-source, single-channel systems, we provide an optimal scheduling policy. In multi-source, multi-channel systems, the scheduling problem becomes a multi-action restless multi-armed bandit problem. For this setting, we design a new scheduling policy by integrating Whittle index-based source selection and duality-based feature selection-from-buffer algorithms. This new scheduling policy is proven to be asymptotically optimal. These scheduling results hold for minimizing general AoI functions (monotonic or non-monotonic). Data-driven evaluations demonstrate the significant advantages of our proposed scheduling policies.
@@ I Introduction
A smaller AoI indicates the presence of more recent information at the receiver. While existing AoI research extensively analyzes and optimizes linear and nonlinear functions of AoI [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], there is a lack of clear understanding regarding the value of fresh information in real-time systems. This gap motivates us to seek a more suitable analysis for evaluating the significance of fresh information for systems using it.
Previous studies [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] have assumed that the performance of real-time systems degrades monotonically as the AoI increases. However, our remote inference experiments (see Sec. II-C) demonstrate that this assumption holds in some scenarios (e.g., video prediction), but not in others (e.g., temperature prediction or reaction prediction). For instance, when predicting the temperature for the next hour, the temperature recorded 24 hours ago is more relevant than the temperature recorded 12 hours ago due to daily weather patterns. Therefore, fresher temperature data is not always better for temperature prediction. Similarly, in a reaction prediction system, if a vehicle A initiates braking, nearby vehicles do not react instantly due to the response delay of drivers and braking systems. Therefore, slightly outdated actions of vehicle A can be more relevant for predicting the reactions of nearby vehicles. These observations highlight that fresher data is not always better, emphasizing the need for a theory to accurately assess the value of fresh information.
In this paper, we introduce a new information-theoretic framework for remote inference to analyze when fresher data is better and when it is not. We also study optimal communication system design for remote inference. Most existing studies in the AoI literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] have focused on designing transmission scheduling strategies to minimize monotonic functions of AoI. However, the design of efficient scheduling policies for optimizing general, potentially non-monotonic functions of AoI remains unexplored. To address this gap, we design the first transmission scheduling policies to minimize general functions of AoI, regardless of whether the functions are monotonic or non-monotonic. The contributions of this paper are summarized as follows:
@@ I-A Related Works
The concept of Age of Information (AoI) has attracted significant research interest; see, e.g., [20, 3, 5, 4, 21, 13, 22, 8, 7, 23, 24, 6, 9, 10, 11, 12, 14, 15, 16, 25, 26, 27, 28, 29, 30] and a recent survey [31]. Initially, research efforts were centered on analyzing and optimizing average AoI and peak AoI in communication networks [3, 5, 4, 21, 13]. Recent research endeavors on AoI have shifted towards optimizing the performance of real-time applications, such as remote estimation [9, 25, 29], remote inference [20], and control systems [11, 30], by leveraging AoI as a tool. In [22, 8, 7, 24], information-theoretic metrics such as Shannon’s mutual information (or Shannon’s conditional entropy) has been used to quantify the amount of information carried by the received data about the current source value (or the amount of uncertainty regarding the current source value) as the data ages. In addition, a Shannon’s conditional entropy term $H_{\mathrm{Shannon}}(Y_{t}|X_{t-\Delta(t)}=x,\Delta(t)=\delta)$ was used in [23] to quantify information uncertainty given the most recent observation $X_{t-\Delta(t)}=x$ and AoI $\Delta(t)=\delta$. The information-theoretic metrics in these prior studies [22, 8, 7, 24, 23] cannot be directly used to evaluate real-world system performance. To bridge the gap, in the present paper, we use an $L$-conditional entropy $H_{L}(Y_{t}|X_{t-\Delta(t)},\Delta(t))$, to approximate and analyze the inference error in remote inference, as well as the estimation error in remote estimation. For example, when the loss function $L(y,\hat{y})$ is chosen as a quadratic function $(y-\hat{y})^{2}$, the $L$-conditional entropy $H_{L}(Y_{t}|X_{t-\Delta(t)},\Delta(t))=\mathbb{E}[(Y_{t}-\mathbb{E}[Y_{t}|X_{t-\Delta(t)},\Delta(t)])^{2}]$ is exactly the minimum mean squared estimation error in signal-agnostic remote estimation. This approach takes a significant step to bridge the gap between AoI metrics and real-world applications, by directly mapping AoI to the application performance metrics.
In earlier AoI studies [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], it was usually assumed that the observed data sequence is Markovian and the performance degradation caused by information aging was modeled as a monotonic AoI function. Hence, the studies [4, 6, 5, 7, 12, 8, 9] adopted the “generate-at-will” status updating model, where the transmitter can only send the most recently generated signal value. However, practical data sequence may not be Markovian [7]. In the present paper, we propose a new local information geometric approach to analyze both Markovian and non-Markovian time-series data. For non-Markovian data, fresher data is not always better. To that end, we propose a new status updating model called the “selection-from-buffer” model, where the transmitter has the option to send any of the $B$ most recent features stored in the source’s buffer.
@@ II-B Offline Training and Online Inference
A supervised learning algorithm consists of two phases: offline training and online inference. In the offline training phase, a neural network is trained using one of the following two approaches.
@@
In the online inference phase, the pre-trained neural network is used to predict the target $Y_{t}$ in real-time. We assume that the process $\{(Y_{t},X_{t}),t=0,1,2,\ldots\}$ is stationary and the processes $\{(Y_{t},X_{t}),t=0,1,2,\ldots\}$ and $\{\Delta(t),t=0,1,2,\ldots\}$ are independent. Under these assumptions, if $\Delta(t)=\delta$, the inference error at time $t$ can be expressed as a function of the AoI value $\delta$, i.e.,
@@
where $P_{Y_{t},X_{t-\delta}}$ is the distribution of the target $Y_{t}$ and the feature $X_{t-\delta}$, and $\phi^{*}$ is the trained neural predictor. The proof of (4) is provided in Appendix E. In Sections V-VI, to minimize inference error, we will develop signal-agnostic transmission scheduling policies in which scheduling decisions are determined without using the knowledge of the signal value of the observed process. If the transmission schedule is signal-agnostic, then $\{(Y_{t},X_{t}),t=0,1,2,\ldots\}$ is independent of the AoI process $\{\Delta(t),t=0,1,2,\ldots\}$.
@@ II-C Experimental Results on Information Freshness
We conduct five remote inference experiments to examine how the training error and the inference error vary as the AoI increases. These experiments include (i) video prediction, (ii) robot state prediction in a leader-follower robotic system, (iii) actuator state prediction under mechanical response delay, (iv) temperature prediction, and (v) wireless channel state information prediction. In these experiments, we consider the quadratic loss function $L(\mathbf{y},\hat{\mathbf{y}})=\|\mathbf{y}-\hat{\mathbf{y}}\|^{2}_{2}$. Detailed settings of these experiments can be found in Appendix A. We present the experimental results of the first training method in Figs. 1-5. Related codes and datasets are accessible in our GitHub repository. [^1] https://github.com/Kamran0153/Impact-of-Data-Freshness-in-Learning To illustrate the training error of the second training method as a function of the AoI $\delta$, one can simply assess the training error using the training data samples with the AoI value $\delta$. When the neural network is sufficiently wide and deep, the results of the two training methods are similar. Hence, the experimental results of the second training method are omitted.
@@ III-A1 Training Error of the First Training Approach
$H_{L}(Y|X)=\min_{\phi(x,\delta)\in\mathcal{A},~{}\!\forall x\in\mathcal{X}}\mathbb{E}_{Y,X\sim P_{Y,X}}[L(Y,\phi(X,\delta))].$ (6)
@@ III-A2 Training Error of the Second Training Approach
$H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\Theta},\Theta)$
$=$ $\min_{\phi\in\Phi}\mathbb{E}_{Y,X,\Theta\sim P_{\tilde{Y}_{0},\tilde{X}_{-\Theta},\Theta}}\![L(Y,\!\phi(X,\Theta))],$ (10)
@@
which connects the information-theoretic lower bounds of $\mathrm{err_{training,1}}(\delta)$ and $\mathrm{err_{training,2}}$.
@@ III-A3 Inference Error
$H_{L}(P_{Y};P_{\tilde{Y}}|P_{X})$
$=$ $\sum_{x\in\mathcal{X}}P_{X}(x)\mathbb{E}_{Y\sim P_{Y|X=x}}\!\!\left[L\!\left(Y,a_{P_{\tilde{Y}|\tilde{X}=x}}\right)\right]\!.$ (15)
If the function spaces $\Lambda$ and $\Phi$ are close to each other, the difference between $\mathrm{err_{inference}}(\delta)$ and the $L$-conditional cross entropy $H_{L}(P_{Y_{t}|X_{t-\delta}};P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}|P_{X_{t-\delta}})$ is small.
@@
We first analyze the monotonocity of the $L$-conditional entropy $H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\delta})$ as $\delta$ increases. If $\tilde{Y}_{0}\leftrightarrow\tilde{X}_{-\mu}\leftrightarrow\tilde{X}_{-\mu-\nu}$ is a Markov chain for all $\mu,\nu\geq 0$, by the data processing inequality for $L$-conditional entropy [38, Lemma 12.1], $H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\delta})$ is a non-decreasing function of $\delta$. Nevertheless, the experimental results in Figs. 1-5 show that the training error is a growing function of the AoI $\delta$ in some systems (see Fig. 1), whereas it is a non-monotonic function of $\delta$ in other systems (see Figs. 2-5). As we will explain below, a fundamental reason behind these phenomena is that practical time-series data for remote inference could be either Markovian or non-Markovian. For non-Markovian $(\tilde{Y}_{0},\tilde{X}_{-\mu},\tilde{X}_{-\mu-\nu})$, $H_{L}(\tilde{Y}_{0}|\tilde{X}_{-\delta})$ is not necessarily monotonic in $\delta$.
@@
A Markov chain is an $\epsilon$-Markov chain with $\epsilon=0$. If $Z\rightarrow X\rightarrow Y$ is a Markov chain, then $Y\rightarrow X\rightarrow Z$ is also a Markov chain [45, p. 34]. A similar property holds for the $\epsilon$-Markov chain.
@@
Figure 5: Performance of channel state information prediction. The training error and inference error are non-monotonic in AoI. As the feature sequence length $u$ increases, the errors tend closer to non-decreasing functions of the AoI.
@@
Using (8), (III-A1), and (III-A3), it is easy to show that the $L$-conditional cross entropy $H_{L}(P_{Y_{t}|X_{t-\delta}};P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}|P_{X_{t-\delta}})$ is lower bounded by the $L$-conditional entropy $H_{L}(Y_{t}|X_{t-\delta})$. In addition, $H_{L}(P_{Y_{t}|X_{t-\delta}};P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}|P_{X_{t-\delta}})$ is close to its lower bound $H_{L}(Y_{t}|X_{t-\delta})$, if the conditional distributions $P_{Y_{t}|X_{t-\delta}}$ and $P_{\tilde{Y}_{0}|\tilde{X}_{-\delta}}$ are close to each other, as stated in Lemma 3.
@@
Given $\beta\geq 0$, if
@@ III-D Interpretation of the Experimental Results
We use Theorems 1-3 to interpret the experimental results in Figs. 1-5. In Fig. 1, the training and inference errors for video prediction are increasing functions of the AoI. This observation suggests that the target and feature time-series data $(Y_{t},X_{t-\mu},X_{t-\mu-\nu})$ for video prediction is close to a Markov chain. As discussed in Appendix A, in the robot state prediction experiment, communication delays and periodic updates from leader robot introduce $20$ to $40$ time slots of AoI in the follower robot’s perception of the leader robot’s state. Consequently, the follower robot receives delayed state information from the leader robot to determine its movement. This implies that slightly outdated state of the leader robot $X_{t-d}$ are more relevant that the fresh observation $X_{t}$ for predicting the current state of the follower robot $Y_{t}$, as illustrated in Fig. 2 and the target and feature data sequence $(Y_{t},X_{t-\mu},X_{t-\mu-\nu})$ may deviate significantly from a Markov chain. In the experiment of actuator state prediction under mechanical response delay, pole angle at time $t$ is strongly correlated with the cart velocity generated $25$ ms ago, as observed from data traces in Fig. 3(b). Moreover, temperature and CSI signals have temporal dependence. For example, temperature at time $t$ depends on the temperature of $24$ hours ago. These observations imply that the target and feature data sequence $(Y_{t},X_{t-\mu},X_{t-\mu-\nu})$ for all $\mu,\nu\geq 0$ may not be close to a Markov chain in the experimental results depicted in Figs. 2-5. Because the target and feature time-series data involved is non-Markovian, Theorems 1-3 suggest that the training error and inference error could be non-monotonic with respect to AoI, as observed in Figs. 2-5.